# Neural Radiance Fields in Medical Imaging: Challenges and Next Steps

### v3

@@ 3.2 Limitations
Although impressive progress has been made in the performance of NeRF in medical images, there are several concerns over the current methods.
Dataset Quality and Diversity. NeRFs require a substantial amount of paired data to train, specifically diverse and high-quality images from various angles and perspectives, see Section 4 for more details. In medical imaging, obtaining such paired datasets can be difficult due to privacy concerns, the rarity of certain conditions, and the technical and ethical challenges associated with collecting and sharing patient data.
Limited Resolution and Detail. While NeRFs can produce impressive 3D reconstructions, capturing the intricate details necessary for medical diagnosis can be challenging. The resolution and fidelity of NeRF-generated images may not yet match the standards required to visualize the fine anatomical structures or pathological changes critical for accurate medical assessments [19].
Handling of Transparent and Reflective Surfaces. Medical images often contain structures that are semi-transparent or have varying degrees of reflectivity, such as soft tissues versus bones. NeRFs, originally designed for opaque objects in natural light settings, may struggle to accurately model these properties, leading to less accurate reconstructions.
@@
Generalization Across Different Imaging Modalities. Medical imaging encompasses a wide range of modalities (e.g., MRI, CT, and X-ray), each with its own unique characteristics. NeRFs developed for one type of imaging may not generalize well on another without significant adjustments or retraining, limiting their versatility [20].
Interpretability and Clinical Validation. For NeRFs to be adopted in clinical practice, their outputs must be interpretable by medical professionals and validated against established medical imaging standards [21]. The black-box nature of deep learning models, including NeRFs, can make it challenging to understand how they generate their outputs, complicating clinical validation and trust.
Dataset Year Modality Organ Method Link
MedNeRF [5] 2022 CT Lung, Knee DRR Github [22]
NAF [17] 2022 CBCT Chest, Jaw, Foot, and Abdomen TIGRE Toolbox Github [23]
PerX2CT Dataset [24] 2023 CT Lung DRR Github [25]
DIF-Net Dataset [26] 2023 CBCT Knee DRR Github [27]
Table 1: Overview of Representative Publicly Available Dataset for Medical Imaging NeRFs.
@@ 4 Datasets and Evaluation Metrics
Public datasets play a pivotal role in the advancement of Neural Radiance Fields (NeRFs) in medical imaging, offering a wealth of diverse and annotated data essential for the development and validation of algorithms.
@@
4.1 Overview
Training Medical Image NeRFs needs datasets comprising paired 2D (e.g. X-ray) and 3D (e.g. CT, MRI) scans. However, assembling a real dataset of such pairs is virtually impractical due to the rarity of obtaining X-ray pairs and CT (or MRI) scans captured simultaneously in clinical practice [24]. For example, gathering paired X-ray and CT data can introduce inaccuracies due to patient movement and variations in equipment. Additionally, this process may expose patients to increased radiation levels [12].
Thus, the Reconstructed Radiographs (DRR) technique is employed for simulating and generating X-ray images [5]. This method entails the creation of virtual X-ray projections using three-dimensional medical image data, such as CT or MRI. By generating DRRs, we can avoid the need for paired X-rays and corresponding CT reconstructions, which would otherwise subject patients to increased radiation exposure. The overview relation between Digitally Reconstructed Radiographs (DRR) and medical imaging NeRFs is shown in Figure 8.
4.2 Digitally Reconstructed Radiographs (DRR)
Digitally Reconstructed Radiographs (DRRs) are synthetic X-ray images generated from three-dimensional (3D) computed tomography (CT) data using digital image processing techniques. This technology allows for the creation of X-ray-like images without the need for additional radiation exposure to the patient. DRRs are produced by simulating the path of X-rays through a 3D volume dataset, mimicking the attenuation and absorption properties of X-rays as they pass through different tissues and structures within the body.
The primary utility of DRRs lies in their application within various medical fields, including radiation oncology, orthopedics, and interventional radiology, among others. In radiation therapy planning, for instance, DRRs are used to accurately target treatment areas, allowing for precise alignment of the therapeutic radiation beam with the tumor while sparing surrounding healthy tissue. Similarly, in orthopedics, DRRs assist in pre-surgical planning and the alignment of prosthetic implants by providing clear images of bone structures.
Furthermore, DRR technology facilitates the comparison and verification of patient positioning during treatment, ensuring that the actual patient alignment matches the planned treatment setup. By leveraging existing CT datasets to generate these radiograph-like images, DRRs eliminate the need for additional diagnostic X-ray examinations, thereby reducing the cumulative radiation dose to the patient and enhancing the efficiency of medical imaging workflows. More DRR generation examples can be found in [28].
@@ 4.3 Public Dataset
MedNeRF Dataset [5] (Github [22]). The dataset includes 20 CT chest scans as in [29] and [30], in addition to 5 CT knee scans documented in [31] and [32]. These scans represented a diverse array of patients, encompassing various contrast modalities and encapsulating both typical and atypical anatomical structures. The DRR generation process involves the emulation of the rotation of a radiation source and imaging panel around the vertical axis, resulting in the creation of DRRs with a resolution of 128 $\times$× 128 at five-degree intervals. Consequently, each object produced a total of 72 distinct DRRs.
@@
PerX2CT Dataset [24] (Github [25]). The dataset is built based on the LIDC-IDRI dataset, the DRR technology as a solution, enabling the generation of synthetic X-rays from actual CT volumes to simulate the required paired data for studies, the dataset is constructed following [33].
DIF-Net Dataset [26] (Github [27]). The dataset comprises a knee CBCT collection with 614 CT scans, distributed as 464 for training, 50 for validation, and 100 for testing purposes. It also includes 2D projections created using DRR technology, each with a resolution of 256 $\times$× 256.
@@ 4.4 Evaluation Metrics
3D Projection Reconstruction. To assess the efficacy of the model in reconstructing CT projections, two principal image quality metrics: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) are used [5]. PSNR focuses on the accuracy of luminance, quantifying the variance between the original and the model-generated images. Meanwhile, SSIM offers a holistic evaluation, taking into account factors such as luminance, contrast, and structural integrity.
Utilizing these metrics is essential for a precise evaluation of the model’s performance in CT projection reconstruction, providing a reliable basis for the research outcomes. These metrics are widely recognized for their effectiveness in measuring the quality of image reconstructions and CT projections, underscoring their importance in the analysis.
@@ 5 FUTURE DIRECTIONS
Besides continuing to address the aforementioned limitations, we also envision a few important directions of NeRFs in medical imaging that will receive more attention in the future.
Enhanced Detail Resolution. Future developments should focus on improving the resolution and detail that NeRFs can capture, especially for internal structures within the medical images. Advanced algorithms could be designed to more effectively learn fine-grained details from limited or sparse imaging data, potentially through the integration of higher-resolution imaging techniques or the development of more sophisticated methods.
@@
Accurate Color and Density Representation. Since color density variations convey vital information in medical images, future NeRF models could benefit from enhancements in accurately modeling and reproducing these variations. This could involve adapting NeRFs to more precisely interpret and generate images based on the organ properties and physiological conditions that these color densities represent, perhaps by integrating additional data sources or modalities that provide complementary information.
@@
Reducing Computational Demands. The computational intensity of training and deploying NeRF models, especially for high-resolution medical imaging data, poses a significant challenge. Future research should aim to optimize these models for greater efficiency, potentially through algorithmic improvements, the use of more efficient neural network architectures, or leveraging hardware accelerations.
Customization and Personalization. Tailoring NeRFs to individual patient data for personalized medicine could greatly enhance diagnostic accuracy and treatment planning. Future directions might include developing adaptable NeRF models that can quickly learn from new patient data or incorporating dynamic adjustments based on real-time imaging during surgical procedures.
@@
Integration with Diagnostic. Exploring the integration of NeRF-generated 3D models into diagnostic workflows, surgical planning, and even real-time image guidance during procedures could have profound implications for patient care. This includes the development of tools and interfaces that allow clinicians to interact with and manipulate 3D reconstructions for better planning and execution of treatments.
Integration with New Advanced Techniques. The future of NeRFs in medical imaging is poised for significant advancements, particularly with the integration of new techniques.
@@
For example, NeRFs could integrate with 3D Gaussian Splatting [37] for real-time radiance field rendering, which promises to address some of the current limitations of NeRFs, such as the high computational cost and the need for faster rendering times without compromising the quality of the 3D reconstructions. 3D Gaussian Splatting, by efficiently approximating the continuous radiance fields with discrete samples, can greatly accelerate the rendering process. This method, when applied to NeRFs in medical imaging, could facilitate real-time visualization of complex anatomical structures, making it invaluable for surgical planning and intraoperative navigation. The computational intensity of NeRFs currently limits their application in scenarios where quick decision-making is critical, such as emergency medicine or real-time surgical assistance. Incorporating 3D Gaussian Splatting could substantially reduce the computational load, enabling the deployment of NeRF-based applications on more accessible hardware, including mobile devices used in remote diagnostics.
@@
Besides, the integration of NeRFs with Reinforcement Learning (RL) [38] in medical imaging represents a forward-looking approach that can significantly improve the efficiency, accuracy, and personalization of medical imaging and related procedures. By learning from interactions between data and environments, RL models can optimize NeRFs to meet specific needs of medical imaging, addressing challenges such as scan time, patient comfort, and the complexity of interpreting medical images [39]. This synergy has potential to transform medical diagnostics and treatment planning, making healthcare more efficient and effective for patients.
Furthermore, as mentioned before, medical images often come with noise and artifacts due to various reasons like patient movement, suboptimal equipment settings, or inherent limitations of imaging modalities. Robust machine learning (RML) algorithms are designed to work effectively even in the presence of such noise [40, 41], ensuring consistent diagnostic performance. Generally, the need for RML in NeRFs becomes increasingly critical in high-stakes domains like medical applications, where the inherent risks significantly heighten the importance of such robustness.
@@
Last but not least, the Foundation Models (FMs) [42] offer a promising approach to address some of the unique challenges faced in medical image analysis and reconstruction [43]. The integration of NeRFs with FMs in medical imaging could lead to significant advancements in accuracy, efficiency, and generalizability across different medial imaging modalities and clinical applications. FMs are trained on vast datasets, enabling them to learn a wide range of features and patterns [44]. When applied to medical imaging NeRFs, these models can help overcome the challenge of data variability across different medical conditions and imaging modalities. The depth and complexity of FMs allow them to capture high-level abstractions and intricate details from medical images [45]. When used to generate or refine NeRFs, these models can potentially improve the accuracy of 3D reconstructions, enhance the detection of diseases, and provide more precise anatomical models for simulation purposes.
@@ 6 Conclusion
In conclusion, the exploration of Neural Radiance Fields (NeRF) in medical imaging illuminates a promising pathway toward enhancing diagnostic precision, treatment planning, and early disease detection. This paper has dissected the multifaceted challenges inherent to applying NeRFs to medical image data, such as adapting to the distinct imaging principles of medical modalities, capturing the intricate internal structures necessary for accurate diagnostics, defining object boundaries with precision, and interpreting color density variations critical for medical analysis. Through a comprehensive examination of current methodologies and an insightful discussion on the limitations and potential of NeRFs, we emphasize the necessity for innovative solutions that can navigate these obstacles.
As we venture into future advancements, the integration of NeRFs with emerging technologies and methodologies holds the promise to significantly propel medical imaging forward, emphasizing the critical need for continued research, collaborative efforts, and the development of novel approaches to fully harness the capabilities of NeRFs in revolutionizing medical imaging.
Acknowledgments. Copyrights of the figures belong to their own authors and/or holders (sources are cited in the captions). The author sincerely thanks all readers for taking the time to read the manuscript. If any related work has been overlooked, please do not hesitate to contact us. The feedback will be greatly appreciated and help us considerably improve this work.
Xin Wang is supported by University at Albany Start-up Grant.

### v4

@@ 3.2 Limitations
Although impressive progress has been made in the performance of NeRFs in medical imaging, several concerns remain regarding the current methods.
Dataset Quality and Diversity. NeRFs require a substantial amount of paired data for training, specifically diverse and high-quality images from various angles and perspectives. See Section 4 for more details. In medical imaging, obtaining such paired datasets can be difficult due to privacy concerns, the rarity of certain conditions and the technical and ethical challenges associated with collecting and sharing patient data.
Limited Resolution and Detail. While NeRFs can generate impressive 3D reconstructions, capturing the intricate details necessary for medical diagnosisremains a challenge. The resolution and fidelity of NeRF-generated images may not yet match the standards required to visualize fine anatomical structures or pathological changes crucial for accurate medical assessments [28].
Handling of Transparent and Reflective Surfaces. Medical images often contain structures that are semi-transparent or exhibit varying degrees of reflectivity, such as soft tissues compared to bones. NeRFs, originally designed for opaque objects in natural light settings, may struggle to accurately model these properties, leading to less accurate reconstructions.
@@
Generalization Across Different Imaging Modalities. Medical imaging encompasses a wide range of modalities (e.g., MRI, CT, and X-ray), each with its own unique characteristics. NeRFs developed for one type of imaging may not generalize well to another without significant adjustments or retraining, limiting their versatility [8].
Interpretability and Clinical Validation. For NeRFs to be adopted in clinical practice, their outputs must be interpretable to medical professionals and validated against established medical imaging standards [29]. The black-box nature of deep learning models, including NeRFs, make it challenging to understand how they generate their outputs, complicating clinical validation and trust.
@@ 4 Datasets and Evaluation Metrics
Training Medical Image NeRFs needs datasets comprising paired 2D (e.g. X-ray) and 3D (e.g. CT, MRI) scans. However, assembling a real dataset of such pairs is virtually impractical due to the rarity of obtaining X-ray pairs and CT (or MRI) scans captured simultaneously in clinical practice [30]. For example, gathering paired X-ray and CT data can introduce inaccuracies due to patient movement and variations in equipment. Additionally, this process may expose patients to increased radiation levels [16]. Thus, the Reconstructed Radiographs (DRR) technique is employed for simulating and generating X-ray images [6]. This method entails the creation of virtual X-ray projections using three-dimensional medical image data, such as CT or MRI. By generating DRRs, we can avoid the need for paired X-rays and corresponding CT reconstructions, which would otherwise subject patients to increased radiation exposure. The overview relation between Digitally Reconstructed Radiographs (DRR) and medical imaging NeRFs is shown in Figure 5.
4.1 Digitally Reconstructed Radiographs (DRR)
Digitally Reconstructed Radiographs (DRRs) are synthetic X-ray images generated from three-dimensional (3D) computed tomography (CT) data using digital image processing techniques. This technology allows for the creation of X-ray-like images without the need for additional radiation exposure to the patient. DRRs are produced by simulating the path of X-rays through a 3D volume dataset, mimicking the attenuation and absorption properties of X-rays as they pass through different tissues and structures within the body. The primary utility of DRRs lies in their application within various medical fields, including radiation oncology, orthopedics, and interventional radiology, among others. In radiation therapy planning, for instance, DRRs are used to accurately target treatment areas, allowing for precise alignment of the therapeutic radiation beam with the tumor while sparing surrounding healthy tissue. DRR technology facilitates the comparison and verification of patient positioning during treatment, ensuring that the actual patient alignment matches the planned treatment setup. By leveraging existing CT datasets to generate these radiograph-like images, DRRs eliminate the need for additional diagnostic X-ray examinations, thereby reducing the cumulative radiation dose to the patient and enhancing the efficiency of medical imaging workflows. More DRR generation examples can be found in [31].
@@ 4.2 Public Dataset
MedNeRF Dataset [6] (Github [32]). The dataset includes 20 CT chest scans as referenced in [33] and [34], in addition to 5 CT knee scans documented in [35] and [36]. The DRR generation process involves the emulation of the rotation of a radiation source and imaging panel around the vertical axis, resulting in the creation of DRRs with a resolution of 128 $\times$× 128 at five-degree intervals. Consequently, each object produces a total of 72 distinct DRRs.
@@
PerX2CT Dataset [30] (Github [38]). The dataset is based on the LIDC-IDRI dataset and utilizes DRR technology to generate synthetic X-rays from actual CT volumes, simulating the required paired data for studies. The dataset is constructed following the methodology in[39].
DIF-Net Dataset [40] (Github [41]). The dataset comprises a knee CBCT collection with 614 CT scans, distributed as follows: 464 for training, 50 for validation, and 100 for testing purposes. It also includes 2D projections created using DRR technology, each with a resolution of 256 $\times$× 256.
@@ 4.3 Evaluation Metrics
3D Projection Reconstruction. To assess the efficacy of the model in reconstructing CT projections, two principal image quality metrics: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) are used [6]. PSNR focuses on the accuracy of luminance, quantifying the variance between the original and the model-generated images. Meanwhile, SSIM offers a holistic evaluation, taking into account factors such as luminance, contrast, and structural integrity. Utilizing these metrics is essential for a precise evaluation of the model’s performance in CT projection reconstruction, providing a reliable basis for the research outcomes. These metrics are widely recognized for their effectiveness in measuring the quality of image reconstructions and CT projections, underscoring their importance in the analysis.
@@ 5 FUTURE DIRECTIONS
Besides continuing to address the aforementioned limitations, we also envision several important directions for NeRFs in medical imaging that will receive more attention in the future.
Enhanced Resolution. Future developments should focus on improving the resolution and detail that NeRFs can capture, especially for internal structures within the medical images. Advanced algorithms could be designed to more effectively learn fine-grained details from limited or sparse imaging data, potentially through the integration of higher-resolution imaging techniques or more sophisticated methods.
@@
Accurate Color and Density Representation. Since color and density variations convey vital information in medical images, future NeRF models could benefit from enhancements in accurately modeling and reproducing these variations. This could involve adapting NeRFs to more precisely interpret and generate images based on the organ properties and physiological conditions that these color densities represent, perhaps by integrating additional data sources or modalities that provide complementary information.
@@
Reducing Computational Demands. The computational intensity of training and deploying NeRF models, especially for high-resolution medical imaging data, poses a significant challenge [45]. Future research should aim to optimize these models for greater efficiency, potentially through algorithmic improvements, the use of more efficient neural network architectures, or leveraging hardware accelerations [46].
@@
Integration with Diagnosis. Exploring the integration of NeRF-generated 3D models into diagnostic workflows, surgical planning, and even real-time image guidance during procedures could have profound implications for patient care. This could involve developing tools and interfaces that allow clinicians to interact with and manipulate 3D reconstructions, improving treatment planning and execution.
@@
Enhancing Real-Time Performance. NeRFs in medical imaging are poised for significant advancements, particularly with the integration of new techniques. NeRFs could integrate with 3D Gaussian Splatting [47] for real-time radiance field rendering, which promises to address some of the current limitations of NeRFs, such as the high computational cost and the need for faster rendering times without compromising the quality of the 3D reconstructions [48, 49]. The computational intensity of NeRFs currently limits their application in time-sensitive scenarios such as emergency medicine or real-time surgical assistance. By incorporating 3D Gaussian Splatting, the computational load could be significantly reduced, making NeRF-based applications more accessible, including mobile devices used in remote diagnostics.
@@
Agentic NeRFs. The integration of NeRFs with Reinforcement Learning (RL) [50] in medical imaging represents a forward-looking approach that can significantly improve the efficiency, accuracy, and personalization of medical imaging and related procedures. By learning from interactions between data and environments, RL models can optimize NeRFs to meet specific needs of medical imaging, addressing challenges such as reducing scan time, improving patient comfort, and simplifying the interpretation of medical images [51].
@@
Foundation Models of NeRFs. Foundation Models (FMs) [52] offer a promising approach to address some of the unique challenges faced in medical image analysis and reconstruction. The integration of NeRFs with FMs in medical imaging could lead to significant advancements in accuracy, efficiency, and generalizability across different medical imaging modalities and clinical applications. FMs are trained on vast datasets, enabling them to learn a wide range of features and patterns [53]. When applied to medical imaging, NeRFs can help overcome the challenge of data variability across different medical conditions and imaging modalities. The depth and complexity of FMs allow them to capture high-level abstractions and intricate details from medical images [54]. When used to generate or refine NeRFs, these models can potentially improve the accuracy of 3D reconstructions, enhance the detection of diseases, and provide more precise anatomical models for simulation purposes.
@@ 6 Conclusion
In conclusion, the exploration of Neural Radiance Fields (NeRF) in medical imaging presents a promising pathway for enhancing diagnostic precision, treatment planning, and early disease detection. This paper has examined the multifaceted challenges of applying NeRFs to medical image data, such as adapting to the distinct imaging principles of medical modalities, capturing intricate internal structures essential for accurate diagnostics, defining object boundaries with precision, and interpreting color and density variations critical for medical analysis. Through a comprehensive examination of current methodologies and an discussion of the limitations and potential of NeRFs, we emphasize the necessity for innovative solutions that can navigate these obstacles.